Science Prep Night 20130208 UT

Summary

Fairly successful night. Clouds in the beginning and one 1-hour fault caused a bit of a rush to finish all the tasks, but overall, I think we were able to confirm that the instruments are ready for science (at least MODS imaging & longslit, LUCI imaging, and LBC). See: https://docs.google.com/a/lbto.org/document/d/1NrDLZ0BsifERWUe_F9CcaGMU-ict4Q5P7NuTyAqE_y8/edit

  • LBC - completed closed dome tests successfully. Ran typical checks (dofpia, rangebal) successfully. While on-image IQ seemed to agree well with the seeing reported by the DIMM, the pupils, especially in blue, did not seem well fit by dofpia around the edges. However, the fairly mediocre (1-1.2") seeing prevented any real exploration of the skirts, other IQ issues etc. Took data for rotcenter, superfoc, and test images for a transit program all successfully. Will need to do analyses ASAP for rotcen and will look into some troubleshooting that I completed for noise and system crashes and update ITs accordingly.

  • MODS - Restarted the ICs but also needed to turn on the lbttcs and imcs. Turns out that the AGW and WFS were also left off. Problems solved and one primary spectrophotometric standard acquired and observed. Instrument seems to be in working order. Single INAF mask should be installed tomorrow.

  • LUCI - closed dome work, including pupil alignment, test of calibration unit, flats, successfully completed. Brief examination on-sky with science target to confirm that imaging mode is indeed working and that AGWs, etc are ok.

LBC (closed dome)

23:45 Started up instrument

23:57 LBC test biases: Count levels ~470-490 Blue and 460-480 Red. Lights on. Data transferring correctly to /newdata

00:14 LBC biases: Count levels ~470-490 Blue and 470-480 Red. Lights off.

00:32 Cracked the vent doors to try to thermalize the mirrors

02:00 We've got significant clouds with rain in the distance and think we can better spend the time starting with LUCI closed dome stuff;

LUCI (closed dome)

01:30 Started up instrument; did a test dark

02:00 Noticed that even after authorization Telescope Control GUI was not updating UT, etc. and getting DD errors. Restarted Telescope server from Engineering GUI

02:05 Reconfiguring

02:22 Take test darks: 2 1x4.0 DCR (0003, 0004); 2 1x12.0 MER 0005, 0006 in dark dome

02:25 Start pupil alignment checks:

  • 0008 inner plug ~ 972; 1080.5. Pupil ~ 966; 1078.5. To move the plug to the cold stop X = -6 and Y = -2
  • 0010 image is +2038, +2044; detector at 2048, 2048. To move image x = +10, y = +4

02:43 Moving FM4:

  • +10 pix in X in image = +1110 steps in Motor 2
  • +4 pix in Y in image = +480 steps in Motor 1

02:50 Image 0012 -- Image is +2044 in X, still need another +4 pix in X

  • +4 pix in X in image = +444 steps in Motor 2

02:53 Image 0014 -- Image is +2046 in Y, still needs another +2 in Y

  • +2 pix in Y in image = +240 steps in Motor 1

02:57 Image 0016 looks great!

03:00 Whoops! realized that M3 wasn't collimated. Collimated and took 0018 to confirm image was still centered on detector. Yes.

03:04 Pupil image:

  • Image 0019 inner plug ~ 890.5; 1171.0 Pupil ~ 890.5, 1161.0; To move the plug to the cold stop X = 0 and Y = -10

FM1; Motor_1: Use: +136steps / +1px = -1360 for y = -10

03:16 Pupil image 0021; looks pretty close!

03:20 Image 0022; looks good. Calling it a win! Pupil and image alignment good smile

03:29 Taking imaging cal with HALO 3

  • 0023-0033: 4s K-band ~ 18000 counts
  • 0034: 4s Ks-band ~ 11000 counts
  • 0035-0044: 5s Ks-band ~ 13000 counts
  • 0045: 10s H-band ~ 26000 counts
  • 0046-0055: 6s H-band ~ 16000 counts
  • 0056: 6s J-band ~ 9000 counts
  • 0057-0066: 10s J-band ~ 16000 counts
  • 0067: 10s z-band ~ 3000 counts
  • 0068-0077: 40s z-band ~ 12000 counts

04:01 Checking HALO1 and HALO2 - Looks good!

  • 4s with Ks and Bry and HALO 1
  • 4s with Ks and Bry and HALO 2

04:05 Reconfiguring for LBC; running flush_detector_x10

09:30 While waiting for MODS issues to be solved, I took the lamp off

  • 0083-0092: 4s LAMP OFF K-band
  • 0093-0102: 5s LAMP OFF Ks-band
  • 0103-0112: 6s LAMP OFF H-band
  • 0113-0122: 10s LAMP OFF J-band
  • 0: 40s LAMP OFF z-band

LBC (on sky)

04:17 Reconfigured but we've got no stars. Waiting it out....

05:40 Finally opening. Slewing to WT10_292 to see if we can collimate, etc.

05:52 Dofpia converges; running co-pointing, successful

06:00 Slewing to Rotcen field; checking dofpia again and testing for IQ. So far it looks like IQ on image matches the DIMM

  • e.g. @6.23 DIMM reporting ~1.1" seeing; V-band image lbcb.20130331.062203.fits = 1.2"; R-band lbcr.20130331.062152.fits = 1.2"

06:24 rotatorcenter_pa0.ob and rotatorcenter_pa180.ob

  • Images are:
    • PA=0; B=062638; PA 180; B=062953
    • PA=0; R=062625; PA 180 R=062949

06:45 running test of guiding and dithering using PG0928+29_2 w/ 200s USpec and 250s z-sloan. All looks good.

07:09 ACT0241 for Superfoc check:

  • dofpia
  • lbcrangebal
  • 7-step superfoc (RB_rVsuperfoc7.ob) at PA=0
  • 7-step superfoc (RB_rVsuperfoc7_pa180.ob) at PA=180
  • dofpia
  • 7-step superfoc (RB_rVsuperfoc7_pa180.ob) at PA=180
  • 7-step superfoc (RB_rVsuperfoc7.ob) at PA=0

08:01 Test of Matthias/Johanna LBC script for transit observation /home/lbto/TestOBs/GJ1214_Teske

  • /home/lbto/TestOBs/GJ1214_Teske
  • execute:
    • dofpia
    • co-pointing script (images: B =080941; R =080936)
    • subframe B = 081222, 081325,081428 ; R = 081216, 081313, 081411
    • full frame B = 081728, 081846, 082003, 082121, 082239, 082357, 082515, 082633, 082750
    • full frame R = 081827, 081938, 082050, 082201, 082313, 082424, 082535, 082647, 082758
  • Separate test - after subframe OB take another full frame OB and then 10 biases, this will check for vertical noise issue
    • Did not see any obvious vertical noise in the images but will have to take a closer look when I analyze the data tomorrow. I did see strong evidence for the type of horizontal news that Jill Gerke and co. were reporting. Again, will need more time off-sky to process and update fault reports

08:40 Quickly do another rotatorcenter_pa0.ob and rotatorcenter_pa180.ob in denser field

  • Images are:
    • PA=0 B= 084219; PA 180; B= 084526
    • PA=0 R= 084226; PA 180 R= 084522

MODS (closed dome and on-sky)

08:50 Restart ICs. Had to also start TCS and IMCS from mods1data

09:07 Sievesnap and general data transfer checks

09:10 First preset, collimation, etc

09:15 Lotso crashes -- see nightlog for details. Taking LUCI Lamp off flats while we wait

10:15 Main problem with secondary solved - however we lost another 10-15 minutes because for reasons I don't understand MODS AGW and WFS were turned off (??) Stephen fortunately knew how to turn them back on from the MODS GUI. Now collimating and moving on quickly

10:25 Primary spectrophotometric standard in all configurations (Hz 44)

LUCI (on-sky)

11:25 Reconfigure for LUCI

11:40 Collimating and WFS check

11:45 Check-out of imaging mode and IQ with guided science observations. DIMM is reading 0.7"; in Ks I'm getting close to 0.55" which is quite good and right on target. Stars look round -- all-in-all a nice data set (0123-0141)

12:39 Calling it a night!

-- MichelleEdwards - 30 Mar 2013
Topic revision: r5 - 31 Mar 2013, MichelleEdwards
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback